Find the maximum value of the function: y=-x^2+8x-4
Found 2 solutions by solver91311, nerdybill:
Answer by solver91311(24713)



Since the lead coefficient is negative, the parabola opens downward. Therefore, the vertex is a maximum. First find the x-coordinate of the vertex. Take the opposite of the coefficient on the first degree term and divide it by 2 times the lead coefficient. The maximum value of the function is the value of the function at the vertex. Take the x-coordinate you just found and evaluate the function at that x-value.


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by nerdybill(7384)

y=-x^2+8x-4
.
Since the coefficient associated with the x^2 is negative -- we know it is a parabola that opens downwards -- the vertex is then the max.
.
value of x at max is:
x = -b/(2a)
x = -8/(2*(-1))
x = -8/(-2)
x = 4
.
plug above back into:
y=-x^2+8x-4
y=-(4)^2+8(4)-4
y=-(16)+32-4
y=-16+32-4
y = 12 (max value of y)